1/4

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Christian Morgenstern,

The Banshee (An Approach)

La lupfantomo The Werewolf Der Werwolf

Varulven

tradukita de Max Knight tradukita de Bertilo Wennergren tradukita de Alexander Gross

tradukita de Sven Collberg

One night, a banshee slunk

away from mate and child, and in

the gloom went to a village teacher’s

tomb, requesting him: “Inflect me,

pray.”

Jen lupfantom’ de fami-

li’ meznokte iam kuris for al tombo de la profesor’. kaj petis pri konjugaci’.

A Werewolf, troubled by his name, Left wife and brood one night and

came To a hidden graveyard to enlist The aid of a long-dead philologist.

Ein Werwolf eines Nachts

entwich von Weib und Kind und sich

begab an eines Dorfschullehrers

Grab und bat ihn: �Bitte, beuge

mich!“ En varulv lopp från fru

m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och

konjugera.

The village teacher climbed up

straight upon his grave stone with its

plate and to the apparition said who meekly knelt before the

dead:

Do diktis tiu dokta ul’ starante sur la kructa-

bul’ dum lup’ senmova kiel

splinto atentis vortojn de l’

mortinto.

“Oh sage, wake up, please don’t be-

rate me,” He howled sadly, “Just conjugate

me.” The seer arose a bit unsteady Yawned twice, wheezed once, and

then was ready.

Der Dorfschulmeister stieg

hinauf auf seines Blechschilds Mes-

singknauf und sprach zum Wolf, der

seine Pfoten geduldig kreuzte vor dem

Toten: mig!” Byskolläraren stod

upp på blecknamnsskyltens mässingknopp

med svar till den, som satt

på pass med tåligt korslagd tass på

tass.

…  …  …  …  …

2/4

“The banSHEE, in the sub-

ject’s place; the banHERS, the possessive

case. The banHER, next, is what

they call objective case–and that is all.”

De “lupfantomo”, simp-

la rekta, tra “lupfintomo”, plej

perfekta, plu “lupfontomo”, iom

rara, ĝis “lupfuntomo”, ho,

koŝmara.

“Well, ‘Werewolf’ is your plural

past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present

tense, And ‘Iswolf,’ ‘Arewolf’ in this same

sense.”

�Der Werwolf“ – sprach der

gute Mann, �des Werwolfs, Genitiv so-

dann, dem Werwolf, Dativ, wie

man’s nennt, den Werwolf, - damit hat’s

ein End!“ “Jag var-ulv 1:a pers. i

sing., du, han, hon, den, det var-

ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingen-

ting.”

The banshee marveled at the

cases and writhed with pleasure,

making faces, but said: “You did not add, so

far, the plural to the singular!”

Ĉi participoj plaĉis

multe al la fantomo, kaj rezul-

te ĝi petis: “Legu plu el

PIV-o pri participoj en pasi-

vo.”

“I know that–I’m no mental

cripple– The future form and participle Are what I crave,” the beast rep-

lied. The scholar paused–again he tried:

Dem Werwolf schmeichelten

die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge

doch zur Einzahl auch die Mehr-

zahl noch!“ Varulven myste glad åt

hur pers. hade följts av sing.

och plur. “Fast”, bad han, “kan det

inte ges ens en enda pers. i tempus pre-

sens?”

The teacher, though, admitted

then that this was not within his

ken. “While bans are frequent”, he

advised, “A she cannot be plurized.”

La viro ĉagreniĝis ta-

men: “Jam fermis mi per fin’

kaj amen, ĉar netransiraj bestoj

feblas kaj la pasiv’ neniam

eblas.”

“A ‘Will-be-wolf?’ It’s just too

long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’

Utterly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

Der Dorfschullehrer aber

musste gestehn, dass er von ihr

nichts wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Si-

nular.

Bekänna måste vår magi-

ster, att därvidlag fanns vissa

brister. En varulv sine imperfecto var contradictio in adjecto.

…  …  …  …  …

3/4

The banshee, rising clammily, wailed: “What about my fami-

ly?” Then, being not a learned

creature, said humbly “Thanks” and left

the teacher.

Sed lupfantomo pri

transiro ja spertas pli ol eĉ vam-

piro. La lup’ revenis larmo-

kula al sia famili’ ulula.

The Werewolf knew better–his sons

still slept At home, and homewards now he

crept, Happy, humble, without apology For such folly of philology.

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib und

Kind! Doch da er keine Gelehrter

eben, so schied er dankend und er-

geben.

Varulven satt med tårad

lins. Han visste att han fanns

och finns. Därtill kom också fru m.m. som även måste existera.

…  …  …  …  …

4/4

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en Esperanton de Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu  la  retejon

http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .